jonny holds his pen on a paper that creates a linear pair. one of the angles measures 13x+7. the measure of the second angle is one less that twice the first angle. find x
Answer by CubeyThePenguin(3113)

measure of first angle = y

y + (2y-1) = 180
3y - 1 = 180
3(13x + 7) = 181
39x + 21 = 181
39x = 170
x = 170/39